Can you please answer the following question:

Delovely Productions Limited wants to raise $40 million for shows. Arights offering will be used to raise the necessary capital from existing shareholders. Each of the eight million shareholders will receive one right. The subscription price is $10 per share and the market price of the existing shares is currently $15.

Calculate the value of a right at the present time.
Answer #1

Value of right = (Stock Price - Subscription price)/(Number of rights for one share+1)

= (15-10)/(2+1)

= $1.6666

i.e. $1.67

Note: Since$40 million are to be raised and number of rights = 8 million and subscription price = $10

Number of rights for one share = 8*10/40 = 2